OpenStudy (anonymous):

if there are 15 heads and 44 legs on a truck how many cows and ducks r on the truck

OpenStudy (the_fizicx99):

There are 7 cows and 8 duck. 7+8 = 15 7*4 = 28 8*2 = 16 28+16 = 44